### v1

# JND-Guided Light-Weight Neural Pre-Filter for Perceptual Image Coding

Abstract
Just Noticeable Distortion (JND)-guided pre-filter is a promising technique for improving the perceptual compression efficiency of image coding. However, existing methods are often computationally expensive, and the field lacks standardized benchmarks for fair comparison. To address these challenges, this paper introduces a twofold contribution. First, we develop and open-source FJNDF-Pytorch, a unified benchmark for Frequency-domain JND-Guided pre-Filters. Second, leveraging this benchmark, we propose a complete learning framework for a novel, lightweight Convolutional Neural Network (CNN). Experimental results demonstrate that our proposed method achieves better compression efficiency, consistently outperforming existing approaches across multiple datasets and encoders. In terms of computational cost, our model is exceptionally lightweight, requiring only 7.15 GFLOPs to process a 1080p image, which is merely 14.1% of the cost of a leading lightweight network in the field. Our work presents a robust solution that demonstrates superior performance and efficiency, supported by a reproducible research platform. The implementation is available at https://github.com/viplab-fudan/FJNDF-Pytorch.
I Introduction
Enhancing the compression efficiency of encoders based on perceptual quality is a core challenge in perceptual coding. Among the various techniques, pre-filter guided by Just Noticeable Distortion (JND) has become a promising approach due to its portability as an encoder-only modification. Existing methods [1, 2, 3, 4, 5, 6, 7, 8, 9] can be broadly categorized by the type of JND model employed: spatial domain [1, 2, 3], frequency domain [4, 5, 6], hybrid spatial-frequency domain [7], and learning-based models [8, 9].
However, current research in this area suffers from two key limitations: 1) the majority of studies rely on Mean Opinion Score (MOS) for video quality evaluation, a metric that is time-consuming and labor-intensive, thereby hindering rapid algorithm iteration and large-scale validation; 2) many works generally fail to employ the Bjøntegaard Delta Bitrate (BD-BR) [10] to evaluate their impact on compression performance, thus lacking an objective quantification of efficiency gains.
While the few studies [5, 6] that utilize BD-BR confirm that traditional frequency-domain methods are effective, they suffer from prohibitive computational complexity. Their reliance on expensive operations like the Discrete Cosine Transform (DCT) and additional detail enhancement modules for salient regions results in a severe increase in encoding time, with overheads reaching up to 313% [5] and 123% [6].
Figure 1: Illustration of our main contributions. (a) We build a standardized benchmark for frequency-domain JND pre-filters. (b) Using this benchmark, we generate a dataset to train our proposed lightweight CNN via a supervised pipeline, which is optimized by a joint spatial and frequency loss function.
In summary, current research on JND-guided pre-filter faces two primary challenges: a limited evaluation methodology, which relies on subjective metrics and lacks standardized benchmarks for objective comparison, and the high complexity of advanced algorithms, which renders the most effective frequency-domain methods impractical for real-time application. This paper addresses these gaps with two main contributions, as illustrated in Fig. 1:
  • We develop and open-source FJNDF-Pytorch, a unified platform for the development and validation of JND pre-filter methods. The platform integrates: 1) mainstream frequency-domain JND modeling and injection methods; 2) multiple standard open-source encoders and datasets; 3) a suite of objective quality metrics, to facilitate reproducible research and address the evaluation gap.
  • We propose a complete framework for lightweight neural pre-filter. Leveraging our benchmark to generate a high-quality dataset, the framework then effectively trains a lightweight network to learn the fundamental characteristics of the reference filter via a supervised, residual learning strategy. The core of our method is a novel hybrid loss function, featuring a frequency-domain loss, which then empowers the network to surpass the performance of this training reference.
II The FJNDF-Pytorch Benchmark
Figure 2: An overview of the FJNDF-Pytorch framework. (a) The modular toolbox, whose components are detailed in Table I, supports two primary pipelines. (b) A benchmarking pipeline for evaluating filters and generating training data. (c) A training pipeline that leverages this data to train learning-based pre-filters.
TABLE I: Components Integrated in the FJNDF-Pytorch Toolbox
Module Integrated Components
JND-Guided Pre-Filter Model Zoo
   Modeling Mainstream DCT-domain models [11, 12, 13, 5]
   Injection Mainstream injection methods [13, 7, 5, 6]
Evaluation & Training Suite
   Dataset Evaluation: HEVC-B [14], XIPH [15], MCL-JCV [16], MCL-JCI [17]
Training: DIV2K [18], KonJND-1K [19]
   Codec x264 [20], x265 [21], libaom [22], VVenC [23] (+ decoders)
Assessment & Optimization Metric
   Metric PSNR, PSNR-HVSM [24], SSIM [25], MS-SSIM [26], VMAF [27], VMAF-NEG [28]
   Loss L1, L2, Charbonnier [29], MS-SSIM [30] Loss
The architecture of our FJNDF-Pytorch framework is illustrated in Fig. 2. Inspired by established platforms such as IQA-Pytorch [31] and BasicSR [32], our benchmark is designed to power a complete research pipeline, from algorithm benchmarking to model training. Due to space constraints, this section focuses on the core components of the JND-guided pre-filter: the Modeling and Injection stages.
II-A JND Modeling
The principle of JND modeling is to quantify the maximum distortion that falls below the perception threshold of Human Visual System (HVS). This is typically achieved via a multiplicative fusion model that estimates a JND threshold $J_{T}$, by integrating key HVS effects, as shown in Eq. (1):
$J_{T}(u,v,N)=s\cdot J_{CSF}\cdot J_{LA}\cdot J_{CM}\cdot J_{SA},$ (1)
where $(u,v)$ are the frequency coordinates within an $N\times N$ DCT block, and $s$ is the summation effect factor. The terms $J_{\text{CSF}}$, $J_{\text{LA}}$, $J_{\text{CM}}$, and $J_{\text{SA}}$ represent the perceptual effects of the Contrast Sensitivity Function (CSF), Luminance Adaptation (LA), Contrast Masking (CM), and Saliency Adaptation (SA).
Our benchmark integrates four representative JND models [11, 12, 13, 5] to cover the field’s breadth and depth. For breadth, we include models [11] and [12], which represent two distinct paradigms for CSF and LA modeling. For depth, we include model [13] (which adds CM to [12]) and model [5] (which adds SA to [13]), forming a clear evolutionary path.
TABLE II: Results of the top-three performing algorithms on VVenC
Dataset Method PSNR PSNR- MS- VMAF- ALL
Ma + Ib HVSM SSIM NEG
HEVC-B [14] [5] + [6] 1.94 -2.83 -1.36 -0.86 -0.78
[13] + [6] 1.95 -2.75 -1.35 -0.67 -0.70
[5] + [7] 4.91 -4.29 -2.04 -0.85 -0.57
XIPH [15] [12] + [7] -0.27 -4.17 -2.76 -2.87 -2.52
[13] + [6] -0.45 -3.95 -3.14 -2.53 -2.52
[5] + [6] -0.44 -3.98 -3.07 -2.54 -2.51
MCL-JCV [16] [13] + [6] 0.32 -2.57 -2.07 -1.50 -1.46
[5] + [6] 0.31 -2.64 -2.05 -1.43 -1.45
[12] + [7] 0.49 -2.64 -1.35 -1.31 -1.20
MCL-JCI [17] [12] + [13] 8.52 -11.11 -5.34 -4.00 -2.99
[5] + [7] 8.84 -10.39 -4.63 -4.85 -2.76
[13] + [7] 8.79 -10.29 -4.46 -4.54 -2.63
aM and bI denotes the method for JND Modeling and Injection, respectively
II-B JND Injection
JND injection translates the computed threshold $J_{T}$ into signal modifications to remove perceptual redundancy. The integrated methods can be categorized into two main strategies. The first, Coefficient Suppression, directly reduces the magnitude of DCT coefficients ($C_{o}$). This includes a range of methods, from basic suppression [13] to more sophisticated approaches. Frequency-weighted suppression [7] introduces a suppression weight $p$ that is dependent on $(u,v)$, and  [6] further condition this weight on the block type $B$. This progressive refinement is generalized by the formula in Eq. (2):
$C_{f}=\begin{cases}sgn(C_{o})\cdot\sqrt{C_{o}^{2}-p(u,v,B)\cdot J_{T}^{2}},&else\\ 0,&if\kern 5.0pt|C_{o}|<J_{T}\end{cases}$ (2)
The second strategy, a Filter Approach [5], implements injection through a low-pass Gaussian filter whose variance is controlled by $J_{T}$.
Leveraging this benchmark, we conduct a comprehensive evaluation by testing all combinations of the integrated JND modeling and injection methods. To provide a clear summary of the most effective strategies, we report the top-three performing combinations for each dataset on VVenC in Table II.
III JND-Guided Lightweight Network
Figure 3: Training pipeline of our lightweight pre-filter network. Our core contribution, the hybrid loss, supervises the network. The spatial-domain loss ($L_{c},L_{m}$) is computed against a Ground Truth (GT) from a reference filter. The frequency-domain loss imposes a residual constraint ($L^{\text{res}}_{\text{dct}}$) against the GT and a conservation constraint ($L^{\text{cons}}_{\text{dct}}$) against the original input. The network backbone (MBR, FST and HDPA) is detailed in [33]. MBR is re-parameterized into standard convolutions at inference pipeline.
Our proposed method is a complete learning framework designed to train a lightweight network for JND-guided pre-filter. As illustrated in Fig. 3, we employ a supervised learning paradigm with a residual learning strategy, which reduces the task’s complexity and makes it feasible for a lightweight network. The framework is composed of three key components: a hybrid loss function designed to enable the network to surpass the performance of its reference, a data generation strategy to provide a high-quality training target, and a customized lightweight network architecture for efficient inference.
III-A Hybrid Loss Function Design
To enable the network to surpass the performance of the traditional reference algorithm, the core of our methodology is a hybrid loss function that incorporates frequency-domain priors. The total loss, ${L}_{all}$, is a weighted sum of a spatial-domain loss and our proposed frequency-domain loss, ${L}_{freq}$:
$L_{all}=\lambda_{1}L_{c}(I_{f},I_{gt})+\lambda_{2}L_{m}(I_{f},I_{gt})+\lambda_{3}L_{freq}(C_{f},C_{gt},C_{o})$ (3)
The spatial-domain loss, guides the network to learn the fundamental behavior of the reference algorithm. As shown in the lower path of Fig. 3, it is computed between the network’s filtered output ($I_{f}$) and the GT ($I_{gt}$) generated by the reference pre-filter. It consists of a Charbonnier [29] loss ($L_{c}$) for pixel-level fidelity and an MS-SSIM [30] loss ($L_{m}$) for multi-scale structural similarity.
Our key innovation is the frequency-domain loss, $L_{freq}$, which combines a reference-dependent distillation term with reference-independent physical constraints, allowing the network to correct the flaws of its training target. It is composed of two terms:
$L_{freq}=L_{dct}^{res}+L_{dct}^{cons}$ (4)
The first term, the Frequency-Domain Residual Loss ($L_{dct}^{res}$), serves as a distillation constraint. It penalizes the energy of the residual between the DCT coefficients of $I_{f}$ and $I_{gt}$, promoting a deeper learning of the reference’s behavior in the frequency domain and is formulated as:
${L}_{dct}^{res}=\sum_{u,v}\left(C_{f}-C_{gt}\right)^{2}$ (5)
The second term is the Frequency Conservation Constraint Loss ($L_{dct}^{cons}$), which references the original input ($I_{o}$) to enforce physical plausibility. As illustrated in Fig. 4, this loss partitions the 8x8 DCT block into Low-Frequency ($LF$) and High-Frequency ($HF$) areas. It then imposes a bidirectional, asymmetric constraint:
$L^{\text{cons}}_{\text{dct}}=\sum_{(u,v)\in LF}\max(0,|C_{o}|-|C_{f}|)^{2}+\sum_{(u,v)\in HF}\max(0,|C_{f}|-|C_{o}|)^{2}$ (6)
The $LF$ term prevents over-smoothing by penalizing the loss of low-frequency energy, while the $HF$ term suppresses artifacts by penalizing the creation of new high-frequency energy. This constraint empowers the network to surpass the reference by correcting its frequency-domain imperfections.
Figure 4: Illustration of the frequency area partitioning for our proposed conservation constraint loss, $L^{\text{cons}}_{\text{dct}}$. An 8x8 DCT block is divided into a Low-Frequency ($LF$) and a High-Frequency ($HF$) area based on a zigzag scan order and a predefined threshold, $K$.
III-B Data Generation Strategy
To generate the training data, we adopte a data-driven approach. Leveraging our benchmark, we select the top-performing traditional pre-filter, [13][6], from Table II as the reference algorithm to process the DIV2K [18] and KonJND-1K [19] datasets, creating 2,008 original-reference image pairs, providing a validated, high-quality learning target for subsequent network training.
III-C Lightweight Network Architecture
We select the efficient MobileNetIE [33] as our backbone, leveraging its re-parameterization mechanism for low inference complexity. To further specialize this general-purpose model for our JND-Guided pre-filter task, we introduce two key architectural modifications. First, we replace all 5x5 convolutions with 3x3 to better align the network’s receptive field with the local nature of DCT-domain operations. Second, we design a bottleneck-expansion channel configuration, as illustrated by the Head, Body, and Tail structure in Fig. 3.
Specifically, while the backbone maintains a relatively uniform channel width across its stages, we narrow the channels in the Head and Body before expanding them in the Tail. This structure encourages the network to learn a compact and refined feature representation in the early layers, enhancing parameter efficiency. These customizations yield an 41% reduction in computational cost over the backbone, resulting in a highly specialized and efficient model.
IV Experiments
TABLE III: Performance comparison of different pre-filter methods across various datasets on VVenC.
Dataset PSNR PSNR-HVSM SSIM MS-SSIM VMAF VMAF-NEG ALL
[5] [6] [9] Ours [5] [6] [9] Ours [5] [6] [9] Ours [5] [6] [9] Ours [5] [6] [9] Ours [5] [6] [9] Ours [5] [6] [9] Ours
HEVC-B 1.55 1.95 1.51 0.87 -1.81 -2.75 -2.53 -4.20 1.84 -0.73 -1.78 -2.05 -0.79 -1.35 -2.34 -2.90 1.50 -0.54 -2.42 -4.84 0.03 -0.67 -1.81 -3.67 0.39 -0.68 -1.56 -2.80
XIPH -0.77 -0.45 -0.31 -1.16 -3.15 -3.95 -3.07 -5.32 0.82 -2.15 -1.98 -2.78 -1.66 -3.14 -2.78 -4.13 0.09 -2.55 -4.12 -6.26 -1.31 -2.53 -3.27 -5.07 -1.00 -2.46 -2.59 -4.12
MCL-JCV 0.45 0.32 0.21 -0.26 -2.00 -2.57 -1.94 -3.64 1.02 -1.21 -1.30 -1.73 -1.33 -2.07 -1.88 -2.69 1.58 -1.38 -2.72 -5.05 0.21 -1.50 -2.12 -3.61 -0.01 -1.40 -1.63 -2.83
MCL-JCI 2.81 4.81 3.73 2.74 -6.49 -6.31 -7.51 -8.80 3.01 -0.17 -2.04 -2.20 -3.20 -2.77 -5.57 -5.98 -0.91 -2.58 -5.62 -7.55 -2.66 -2.77 -4.88 -6.55 -1.24 -1.63 -3.65 -4.72
TABLE IV: Performance comparison of different pre-filter methods across various encoders on HEVC-B.
Encoder Method PSNR PSNR- MS- VMAF- ALL
HVSM SSIM NEG
libaom Ours -0.34 -5.70 -4.71 -6.20 -4.24
 [9] 0.67 -3.26 -3.49 -3.29 -2.34
[6] 1.42 -3.25 -2.14 -1.48 -1.36
[5] 1.30 -2.43 -1.26 0.66 -0.43
x265 Ours 0.00 -6.81 -4.47 -5.66 -4.23
 [9] 1.01 -3.85 -3.31 -2.67 -2.21
[6] 2.04 -5.15 -2.69 -1.48 -1.82
[5] 0.58 -3.98 -2.44 0.05 -1.45
x264 Ours -1.51 -7.17 -5.93 -5.60 -5.05
 [9] 0.10 -4.06 -4.31 -3.00 -2.82
[6] -0.52 -6.55 -4.86 -2.45 -3.60
[5] -2.00 -5.93 -4.90 -1.55 -3.59
IV-A Experimental Setup
All experiments are conducted using our FJNDF-Pytorch benchmark under the All-Intra configuration with QP values {27, 32, 37, 42}, measuring BD-BR [10] savings (%). We evaluate our method against three state-of-the-art pre-filters [5, 6, 9] on four encoders [20, 21, 22, 23] and datasets [14, 15, 16, 17], as detailed in Table I. In line with the competing methods, we process only the luma channel in all experiments.
For training, we use the Adam optimizer for 100K iterations with a batch size of 16 on 224x224 patches and an initial learning rate of 3e-4. The loss weights ($\lambda_{1}$$\lambda_{2}$ and $\lambda_{3}$) for $L_{c}$, $L_{m}$, and $L_{freq}$ are set to 1.0, 0.16, and 0.02, respectively, with a zigzag cutoff $K=10$ for $L^{cons}_{dct}$.
IV-B Experimental Results
As shown in Table III, our method consistently outperforms both state-of-the-art competitors on the VVenC encoder across all datasets. For instance, on the HEVC-B dataset, our approach achieves a remarkable -4.84% BD-BR saving in VMAF. Furthermore, our approach demonstrates its robustness by securing the highest overall performance on all four datasets, with BD-BR savings ranging from -2.80% to -4.72%.
To validate its generalizability, we conduct further comparisons on HEVC-B across libaom, x265, and x264 encoders. Table IV confirm that our method consistently achieves the best overall performance across all tested encoders, with significant gains of -4.24%, -4.23%, and -5.05%, respectively. This highlights the robustness and wide applicability of our approach.
Due to space constraints, we cannot present the full results for all combinations of datasets and encoders in this paper. However, these comprehensive results are documented and available in our benchmark for reference.
IV-C Ablation Study
We conduct ablation studies to validate our key contributions, with results in Table V. The Baseline model, trained with only spatial-domain loss, surpasses the Reference in perceptual metrics but degrades PSNR. The introduction of our proposed loss, $L_{freq}$, not only rectifies the PSNR degradation, but also boosts perceptual performance in VMAF, confirming the effectiveness of our frequency-domain loss.
Table V also shows that our final lightweight model, compared to the baseline model, retains the majority of the perceptual optimization capability while operating at a significantly lower computational cost. As detailed in Table VI, our final model requires only 7.15 GFLOPs and 1.86K parameters. In addition, we tested the inference time for different methods on both CPU and GPU platform. Methods [5, 6] are our re-implementations, and the measured times are comparable to those reported in [6]. These results validate that our network design is state-of-the-art in both efficiency and effectiveness.
TABLE V: Ablation study of different components of HEVC-B on VVenC.
Method PSNR PSNR-HVSM MS-SSIM VMAF-NEG ALL
Reference $1.95$ 2.75-2.75 1.35-1.35 0.67-0.67 0.70-0.70
BaseLine $3.52$ 5.50-5.50 3.50-3.50 2.85-2.85 2.08-2.08
BaseLine + $L_{freq}$ $0.69$ 4.43-4.43 3.34-3.34 4.18-4.18 2.81-2.81
Ours (Lightweight) $0.87$ 4.20-4.20 2.90-2.90 3.67-3.67 2.47-2.47
TABLE VI: Model size, computational cost, and inference latency.
Method Parameters (K) GFLOPs Inference Time
Train Inference CPUa (s) GPUb (ms)
[5] / / / $\sim 70$ /
[6] / / / $\sim 20$ /
[9] 12.23 12.22 50.70 0.835 15.43
BaseLine 38.92 3.06 12.15 0.595 13.04
Ours 24.95 1.86 7.15 0.428 10.88
aIntel(R) Xeon(R) Gold 6230 CPU. bNVIDIA GeForce RTX 2080 Ti GPU.
V Conclusion
In this paper, we propose a learning framework for a lightweight JND-guided pre-filter network. The core of our method is a novel frequency-domain loss that references the original input signal to correct the inherent flaws of its training target. Our resulting model achieves a new state-of-the-art balance between compression efficiency and computational cost, demonstrating superiority across multiple encoders. The open-sourced FJNDF-Pytorch platform provides a solid foundation for these contributions and for future research in this area. Building on this work, we plan to expand our benchmark to comprehensively evaluate a wider array of methods, including spatial, temporal, and other learning-based approaches.